# Assessing (im)balance in signed brain networks

### v1

@@ VIII.1 Consistency checks
Let us start our analysis by verifying that the ensemble distribution of each, pair-specific signature is consistent with its analytical counterpart, for all benchmarks considered here.
[…]
@@ VIII.2.1 Connectance
When considering the unsigned case, the connectance of the projection returned by the naïve approach is typically large Saracco et al. (2017): this holds true here as well, for all the considered subjects (see fig. 4; for a graphical example, see fig. 10). To be noticed that many validated links are positive, an evidence letting us conclude that the time series associated with the related pairs of areas are concordant for the vast majority of the snapshots constituting the considered time window.
@@
Figure 5: Distributions of CV+, CV- and signed degrees. Top panels: box-plots summing up the distributions of the coefficient of variation for positive degrees (CV+, left) and negative degrees (CV-, right) across the three, different projections of our 100100100 subjects. Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: the positive variant of the CV suggests that the corresponding subgraphs are characterised by a quite homogeneous structure while the negative one suggests that the opposite holds true, being the corresponding subgraphs characterised by fat-tailed degree distributions. Middle and bottom panels: distributions of the positive (left) and negative (right) degrees populating our projections, for subjects 106016 (middle panels) and 857263 (bottom panels). Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. While positive distributions appear (approximately) bell-shaped around the average, negative distributions are characterised by fat tails, an evidence suggesting the presence of hubs. Notice that the naïve and bSCM-induced distributions overlap, in the right panels.
@@
The local, bSCM-based filter, instead, accounts for nodes heterogeneity: whenever the positive degree of a node is (much) larger than its negative degree, the distribution of its similarity with any other node moves to the (far) right, hence validating negative links to a much larger extent; the local filter is, thus, stricter with positive than with negative links - and their number is, somehow, re-balanced towards the naïve value.
Figure 6: Distribution of the Jaccard similarity between pairs of projections. The numerator of such an index represents the number of positive (negative) negative links occupying the same position in the adjacency matrix of both subjects and its denominator represents the cardinality of the union of the two sets of positive (negative) links populating the adjacency matrices to be compared. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. Since a value of 111 indicates perfect overlap while a value of 0 indicates no overlap, positive links yield a larger Jaccard similarity than negative links, a result indicating the former ones as more spatially coherent than the latter ones.
@@ VIII.2.2 Tendency to make hubs
A related index is the one named tendency to make hubs (TMH), introduced in Saberi et al. (2021) and quantifying the tendency of a given degree distribution - when dealing with signed graphs, both the positive and the negative one - to host hubs. Hereby, we adopt a slightly different index, rooted in statistical theory and reducing the functional form of a given degree distribution to just one number. Named coefficient of variation (CV), it is defined as the ratio between the standard deviation and the expected value of the reference distribution: in symbols,
@@
As evident from fig. 5, none of our projections seems to be characterised by the presence of positive hubs; when considering the negative degrees, instead, all projections display a right-skewed distribution that suggests the presence of nodes whose number of connections is (much) larger than that of the others: in a sense, the subgraph induced by the positive links tends to be more homogeneous than the subgraph induced by the negative ones [^17] As highlighted in Luppi et al. (2024), ‘[…] whereas cooperative (positive-valued) interactions in the effective connectivity tend to be strong, modular and relatively short-range, competitive interactions are weaker but more long-range, more diffuse and less clustered […]’..
[…]
A related result is the one concerning the degree distributions of cortical and subcortical areas: as evident from the figures in Appendix C, subcortical areas tend to have a larger negative degree than cortical areas.
@@ VIII.2.3 Spatial coherence of signed connections
From what we have learned so far, the positive subgraphs are (overall) structurally homogeneous - positive degree distributions are bell-shaped and well-represented by the average - while the negative ones are (overall) structurally heterogeneous - negative degree distributions are fat-tailed, hence over-dispersed around the average. Let us, now, ask ourselves how similar positive and negative subgraphs are across subjects.
Figure 7: Distributions of the Strong and Weak Degree of Frustration. Box-plots summing up the distributions of the Strong Degree of Frustration (SDoF) and the Weak Degree of Frustration (WDoF) across the three, different projections of our 100100100 subjects. Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: in both cases, the bSRGM-induced projections are characterised by the smallest median and largest concentration while the naïve and bSCM-induced projections are peaked in correspondence of largest values but show more variability. Since $T^{(---)}\simeq 0$, the SDoF and the WDoF are practically identical and both positive, a result confirming the non null level of frustration characterising our brain networks.
In order to answer this question, we calculated the Jaccard similarity between pairs of (adjacency matrices representing) individual projections, per benchmark. As fig. 6 shows, the positive variant of such an index (i.e. the percentage of positive links located in the same position) ranges between 0.60.60.6 and 111 while its negative variant (i.e. the percentage of negative links located in the same position) ranges between 0 and 0.40.40.4: in words, the spatial coherence of positive links is larger than the spatial coherence of negative links, which appear as those (better) characterising single subjects [^19] As highlighted in Luppi et al. (2024), ‘[…] Additionally, we find that the model with competitive interactions is not only better able to reflect the empirical FC that it is intended to model but is also better at distinguishing between the FC patterns of different individuals […]’ and ‘[…] Thus, we conclude that the model with competitive interactions is significantly more individual-specific […]’.
Figure 8: Heatmaps of signed connections within and between brain regions. Values are averaged across subjects, for each of our benchmarks. The intensity of each colour is proportional to the magnitude of the corresponding entry. Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the largest number of positive links is found within the DMN, the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
@@ VIII.2.4 Assessing the brain (im)balance at the microscale
Let us, now, inspect the way negative links are arranged within projections: to this aim, one may consider the index named Strong Degree of Balance (SDoB) Gallo et al. (2024a) and quantifying the percentage of triads characterised by an even number of negative links, i.e. either zero or two - in other words, the ones that are balanced according to the Strong Balance Theory (SBT). Here, instead, we consider its complementary to 111 and name it [^20] A natural choice would have been that of naming it Strong Degree of Frustration (SDoF); since, however, the triad ()(---)( - - - ) is heavily under-expressed (see also fig. 7), the strong and weak versions practically coincide: we have, thus, avoided to explicitly distinguish between the two. Such an evidence has been also highlighted in Luppi et al. (2024), where it is reported that ‘[…] it is less likely that competitively-interacting neighbours of a node will themselves be interacting competitively […]’ Index of Brain Imbalance (IBI): more quantitatively,
$\text{IBI}=\frac{T^{(+-+)}+T^{(---)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}}$ (82)
@@
naturally, the closer the IBI value to 111, the larger the amount of unbalanced triads.
[…]
@@
Let us, now, consider the brain parcellation carried out in Thomas Yeo et al. (2011), where the authors propose a functional segmentation of brain areas by distinguishing regions (hereby, Yeo’s regions) involved in sensory tasks (e.g. the visual and the somatomotor networks) from those involved in higher-order, cognitive tasks (e.g. the DMN): they are named control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the analysis carried out in Demertzi et al. (2022) highlights the presence of negative correlations between the DMN and the so-called task-positive regions, the one carried out in Saberi et al. (2022) shows that, in healthy individuals, unbalanced triads predominantly occur between different Yeo’s regions.
Figure 9: Tensor of the IBI values associated with the triplets of Yeo’s regions. Values are averaged across subjects, for each of our benchmarks. The size of each point is proportional to the standard deviation of the corresponding value. Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the largest IBI values characterise the triplets where the DMN is present once or, at most, twice, the smallest values characterise those belonging to the same Yeo’s region (especially the SM, SVA and VIS networks) - consistently across all benchmarks.
@@
In order to verify the first of the aforementioned findings, we have constructed two $\times 7$ 7 matrices: each entry of the first (second) one encodes the number of positive (negative) links between the corresponding Yeo’s regions, aggregated across individuals. As fig. 8 shows, i) the largest number of positive links is found within the DMN; ii) the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
In order to verify the second of the aforementioned findings, we have constructed a $\times 7\times 7$ 7 tensor, each entry of which encodes the IBI associated with a given triplet of Yeo’s regions, averaged across individuals (see fig. 9). Still, as the triplets with a smaller IBI exhibit a larger inter-subject variability, we have also considered the distribution of such values: according to the MW test, i) the triads belonging to the same Yeo’s region are those exhibiting the lowest level of imbalance, the smallest values characterising the triplets involving nodes from the SM, SVA and VIS networks; ii) the triads where the DMN is present once or, at most, twice, exhibit the highest level of imbalance - consistently across all benchmarks.
@@ VIII.3 Patterns of brain (im)balance at the mesoscale
Following Gallo et al. (2024b), one can adopt an ‘agnostic’ attitude and explore the mesoscale organisation of a projection without aligning with any specific, conceptual framework: a principled approach to achieve such a goal is that of minimising the Bayesian Information Criterion (BIC), i.e.
@@
the first addendum proxies the complexity of a model with the number of its parameters, $\kappa$, the second addendum proxies the accuracy of a model with its log-likelihood, $\ln\mathcal{L}$, and $V=N(N-1)/2$ accounts for the system dimensions.
Figure 10: Community detection for subject $\bm{106016}$. Top panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the Bayesian Information Criterion (BIC). Although the number of detected communities varies from 555 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time. Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the frustration ($F$). The number of detected communities varies from 222 to 444 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.
@@
and a number of parameters $\kappa_{\text{SSBM}}=k(k+1)$, amounting at twice the number of modules, $k$, into which the projection can be partitioned. Naturally, $N_{r}$ is the number of nodes constituting block $r$, $p_{rr}^{+}$ is induced by the empirical number of positive links within block $r$, i.e.
@@
Figure 11: Community detection for subject $\bm{857263}$. Top panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the Bayesian Information Criterion (BIC). Although the number of detected communities varies from 666 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time. Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the frustration ($F$). The number of detected communities is at most 222 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.

### v3

@@ VIII.1 Consistency checks
Let us start our analysis by verifying that the ensemble distribution of each pair-specific signature is consistent with its analytical counterpart, for all benchmarks considered here.
Specifically, let us focus on each subject and test i) if the distribution defined by eq. 44 is compatible with the ensemble distribution of $S_{ij}$, $\forall\>i<j$ induced by the bSRGM; ii) if the distribution defined by eq. 50 is compatible with the ensemble distribution of $S_{ij}$, $\forall\>i<j$ induced by the bSCM. Each subject can be associated with a KS-score, i.e. the percentage of times the Kolmogorov-Smirnov (KS) test (at the level of $5\%$) does not reject the hypothesis of compatibility: while a value of $1$ indicates perfect compatibility, a value of 0 indicates perfect incompatibility. The results are summed up by the box-plots depicted in fig. 3: the median KS-score is $\text{KS}_{5\%}\simeq 0.965$ under both the bSRGM and the bSCM, the $95\%$ of these distributions ranging between the $2.5$-th and the $97.5$-th percentiles reading $q_{2.5},q_{97.5}]=[0.961,0.970]$; in words, our sampling procedure satisfactorily reproduces the analytical distributions defined by eq. 44 and eq. 50 even when the KS test rejects the hypothesis of compatibility (e.g. for subject #100307\#100307).
Figure 5: Distributions of CV+, CV- and signed degrees. Top panels: box-plots summing up the distributions of the coefficient of variation for positive degrees (CV+, left) and negative degrees (CV-, right) across the three different projections of our $100$ subjects. Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion for each model: the positive variant of the CV suggests that the corresponding subgraphs are characterised by a quite homogeneous structure, while the negative one suggests that the opposite holds true; yet, these differences are levelled out when considering bSCM-induced projections. Bottom panels: distributions of the positive (left) and negative (right) degrees populating our projections, pooled across our subjects. While naïve and bSRGM-induced positive distributions appear as left-skewed, their negative counterparts appear as right-skewed; bSCM-induced distributions, instead, are much flatter in both cases, suggesting a larger heterogeneity of the degrees of this kind. The leftmost peak on $k^{+}$ and the rightmost peak on $k^{-}$ are due to subcortical areas.
@@ VIII.2.1 Connectance
When considering the unsigned case, the connectance of the projection returned by the naïve approach is typically large saracco2017: this holds true here as well, for all the considered subjects (see fig. 4). It is worth noting that many validated links are positive. This leads us to conclude that the time series associated with the related pairs of areas are concordant for the vast majority of the snapshots constituting the considered time window.
@@
The local, bSCM-based filter, instead, accounts for nodes heterogeneity: whenever the positive degree of a node is (much) larger than its negative degree, the distribution of its similarity with any other node moves to the (far) right, hence validating negative links to a much larger extent; the local filter is, thus, stricter with positive than with negative links - and their number is, somehow, re-balanced.
Figure 6: Distribution of the Jaccard similarity between pairs of projections. The numerator of such an index represents the number of positive (negative) links occupying the same position in the adjacency matrix of the two subjects to compare and its denominator represents the cardinality of the union of the two sets of positive (negative) links. Bin widths are computed according to the Freedman-Diaconis rule. Averages are plotted as vertical, dashed lines. Since a value of $1$ indicates perfect overlap while a value of 0 indicates no overlap, positive links yield a larger Jaccard similarity than negative links on both the naïve and bSRGM-induced projections, a result indicating the former ones as more spatially coherent than the latter ones. When considering bSCM-induced projections, instead, the spatial coherence of the two kinds of links appears very similar.
@@ VIII.2.2 Tendency to make hubs
A related index is the one named tendency to make hubs (TMH), introduced in Saberi2021 and quantifying the tendency of a given degree distribution - when dealing with signed graphs, both the positive and the negative degree distributions need to be considered - to host hubs. Hereby, we adopt a slightly different index, rooted in statistical theory and reducing the functional form of a given degree distribution to just one number: named coefficient of variation (CV), it is defined as the ratio between the standard deviation and the expected value of the reference distribution, i.e.
@@
while a CV smaller than $1$ indicates that the expected value is representative of the entire distribution (as its dispersion around the average is smaller than it), observing a CV larger than 1 indicates that the expected value is not representative of the entire distribution (as its dispersion around the average is larger than it) - from this perspective, a fat-tailed distribution is expected to obey $\text{CV}>1$.
@@
We also explicitly show the distributions of the positive and negative degrees pooled across our subjects. As KS tests (at the level of $5\%$) reveal, practically all pairwise comparisons point out our three (naïve, bSRGM-induced, and bSCM-induced) positive degree distributions as significantly different; moreover, the appearance of positive hubs is enhanced when the global filter is used.
A related result concerns the degree distributions of cortical and subcortical areas: as evident from the figure in Appendix C, subcortical areas tend to have a larger negative degree than cortical areas.
@@ VIII.2.3 Spatial coherence of signed connections
From what we have learned so far, the positive subgraphs are (overall) structurally homogeneous, while the negative ones are (overall) structurally heterogeneous. Let us, now, ask ourselves how similar positive and negative subgraphs are across subjects.
Figure 7: Distributions of the IBI values across subjects. Box-plots summing up the distributions of the values of the strong (left) and weak (right) variants of the Index of Brain Imbalance (IBI) across the three different projections of our $100$ subjects. Each box-plot proxies the distribution of the corresponding metric, indicating its central tendency and dispersion: in both cases, the bSRGM-induced projections are characterised by the smallest median, while the naïve projections are peaked in correspondence of larger values, although showing more variability. The difference between the ISBI and the IWBI values characterising the bSCM-induced projections is due to the sensitivity of this benchmark to the abundance of the triad ()(---), whose presence is enhanced by the predominantly negative relationships observed within the subcortical and the limbic regions (see also fig. 8). In any case, the positivity of the IBI values confirms the non-null level of frustration characterising humans’ brain networks.
In order to answer this question, we calculated the Jaccard similarity between pairs of (adjacency matrices representing) individual projections, per benchmark. As fig. 6 shows, the positive variant of such an index (i.e. the percentage of positive links located in the same position) ranges between $0.6$ and $1$ while its negative variant (i.e. the percentage of negative links located in the same position) ranges between 0 and $0.4$, on both the naïve and bSRGM-induced projections: in words, the spatial coherence of positive links is larger than the spatial coherence of negative links, which appear as those (better) characterising single subjects. This picture changes when bSCM-induced projections are considered: in such a case, in fact, the spatial coherence of the two kinds of links appears as very similar.
@@ VIII.2.4 Assessing the brain (im)balance at the microscale
Let us now inspect the arrangement of negative links within projections. To this aim, two choices are possible: i) the first one amounts at considering the index named Strong Degree of Balance (SDoB) gallo2024testing and counting the triads characterised by an even number of negative links, i.e. either zero or two - in other words, the ones that are balanced according to the Strong Balance Theory (SBT); ii) the second one amounts at considering the index named Weak Degree of Balance (WDoB) gallo2024testing and counting the triads characterised by zero, two or three negative links - in other words, the ones that are balanced according to the Weak Balance Theory (WBT).
Figure 8: Heatmaps of the signed connections within and between Yeo’s brain regions. Values are averaged across subjects for each of our benchmarks. The intensity of each colour is proportional to the magnitude of the corresponding entry. Network labels stand for subcortical (SUBC), control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks, and default mode network (DMN). The neatest signal is observed for the bSCM-induced projections: while the largest number of positive links is found within the DMN, a large number of negative links is found between the DMN and every other area, as well as between the SUBC and every other area - the largest number of negative connections being located precisely between the DMN and the SUBC. Notice that, concerning the blocks lying on the main diagonal, $L^{-}>L^{+}$ (only) for the SUBC and LIM regions (see also fig. 7).
The indices above induce two variants of the Index of Brain Imbalance (IBI - practically, their complementary to $1$), i.e. the Index of Strong Brain Imbalance
ISBI $=\frac{T^{(+-+)}+T^{(---)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}}$ (82)
and the Index of Weak Brain Imbalance
IWBI $=\frac{T^{(+-+)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}},$ (83)
@@
naturally, the closer their value to $1$, the larger the amount of unbalanced triads.
As fig. 7 shows, the bSRGM-induced projections are the ones characterised by the largest amount of balanced triads - a result that is not surprising once considering the small amount of negative links populating them; the naïve projections, instead, are characterised by a number of balanced triads that amounts at five-to-six times the number of unbalanced triads. Lastly, and most interestingly, the behaviour of the bSCM-induced projections is sensitive to the one of the triad ()(---), that can be quite abundant: such a behaviour characterises the subcortical and the limbic regions, partly reconciling our findings with those in Luppi2024181818Speaking of cortical regions, there it is reported that ‘[…] it is less likely that competitively-interacting neighbours of a node will themselves be interacting competitively […]’.. Overall, however, each median IBI value is positive, a result confirming the non-null level of frustration characterising humans’ brain networks: if, according to Saberi2021, the value $\text{IBI}=0$ is interpreted as the one characterising the energy ground state, the aforementioned results imply that humans’ brain networks are found in some sort of excited state.
Figure 9: Distributions of the ISBI values associated with triplets of brain areas. Cumulative distributions of the ISBI values across the three, different projections of our $100$ subjects: top panels refer to all areas, middle panels refer to all areas without subcortical ones, bottom panels refer to all areas without subcortical and limbic ones. The largest ISBI values characterise the triplets of subcortical regions: removing them leads to a different picture as the largest ISBI values, now, characterise the triplets within the LIM region; removing them as well reveals that the largest ISBI values characterise the triplets where the DMN is present once or twice - and, to a lesser extent, those within the DMN and CONT regions. The smallest ISBI values, instead, characterise those belonging to the same Yeo’s region (especially the DA, SM and SVA networks).
@@
Let us, now, consider the brain parcellation carried out in yeo2011organization, where the authors propose a functional segmentation of brain areas by distinguishing regions (hereby, Yeo’s regions) involved in sensory tasks (e.g. the visual and the somatomotor networks) from those involved in higher-order, cognitive tasks (e.g. the DMN): they are named control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the analysis carried out in Demertzi2022 highlights the presence of negative correlations between the DMN and the other regions, the one carried out in Saberi2022 shows that, in healthy individuals, unbalanced triads predominantly occur between different Yeo’s regions.
@@
In order to verify the first of the aforementioned findings, we have constructed two $8\times 8$ matrices: each entry of the first (second) one encodes the number of positive (negative) links between the corresponding Yeo’s regions - to which the subcortical areas (SUBC) have been added - aggregated across individuals. As fig. 8 shows, i) the largest number of positive links is found within the DMN; ii) a large number of negative links is found between the DMN and every other area, as well as between the SUBC and every other area - quite consistently across benchmarks.
In order to verify the second finding above, we have inspected the distributions of the ISBI values associated with the admissible triplets of areas: as evident from fig. 9, i) the subcortical triplets are those exhibiting the highest level of imbalance [^19] As highlighted in Saberi2022, ‘Our results suggest that subcortical regions have a prominent role in frustration formation in the brain network at both nodal and connectional levels. So they can bring instability and alter properties that facilitate systemic level neural changes, providing adaptive characteristics to the brain’.; ii) the cortical triplets belonging to the LIM region exhibit the highest level of imbalance, followed by those where the DMN is present once or, at most, twice; iii) the cortical triplets belonging to the same Yeo’s region are those exhibiting the lowest level of imbalance (the smallest values characterising the triplets involving nodes from the DA, SM and SVA networks) - with the exception of the LIM region and, to a lesser extent, the DMN and CONT regions.
Figure 10: Signed communities partitioning the brain of subject $\bm{\#106016}$. Top panels: heatmaps illustrating the communities partitioning the projections of subject #106016\#106016, detected by minimising the Bayesian Information Criterion (BIC). Bottom panels: heatmaps illustrating the communities partitioning the projections of subject #106016\#106016, detected by minimising the frustration ($F$). Although the number of detected communities varies, the overall message conveyed by the two methods can be summed up as follows: the brain seems to be constituted by groups of areas working ‘in phase’ most of the time and ‘in counterphase’ with those of other groups; still, BIC minimisation admits areas from the same group working ‘in counterphase’, as well as areas from different groups working ‘in phase’ - while $F$ minimisation does not.
@@ VIII.3 Patterns of brain (im)balance at the mesoscale
Following gallo2024assessing, one can adopt an ‘agnostic’ attitude and explore the mesoscale organisation of a projection without aligning with any specific conceptual framework. A principled approach to achieve such a goal is that of minimising the Bayesian Information Criterion (BIC), i.e.
@@
the first addendum proxies the complexity of a model with the number of its parameters, $\kappa$, the second addendum proxies the accuracy of a model with its log-likelihood, $\ln\mathcal{L}$, and V=N(N1)/2V=N(N-1)/2 accounts for the system dimensions.
@@
and a number of parameters $\kappa_{\text{SSBM}}=k(k+1)$, with $k$ indicating the number of modules into which the projection is partitioned. Naturally, $N_{r}$ is the number of nodes constituting block $r$, $p_{rr}^{+}$ is induced by the empirical number of positive links within block $r$, i.e.
@@
Figure 11: Heatmaps of the communities partitioning the bSCM-induced projections of four subjects. Matrices are coloured according to eq. 97 and eq. 98, i.e. blocks are blue (red) if the majority of links populating them is positive (negative); white blocks on the diagonal indicate singletons and white blocks off the diagonal indicate that $p_{rs}^{+}=p_{rs}^{-}$. Top panels: the communities partitioning the bSCM-induced projections of our subjects are detected by minimising the Bayesian Information Criterion (BIC); all subjects exhibit diagonal blocks with a majority of negative links and off-diagonal blocks with a majority of positive links, thereby satisfying the statistical variant of the Relaxed Balance Theory. Bottom panels: the communities partitioning the bSCM-induced projections of our subjects are detected by minimising the frustration ($F$); (left aside the white ones) all subjects admit blocks with a majority of negative links only off the main diagonal and blocks with a majority of positive links only on the main diagonal, hence aligning better with the statistical variant of the Traditional Balance Theory.